OpenStudy (anonymous):

The midpoint of UV is (5, -11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V.

OpenStudy (anonymous):

@iambatman @uri ?

OpenStudy (anonymous):

@$w3G_Godd

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Ok, take it one component at a time. Let's do x component of the midpoint. Mid point formula for x is \[\frac{x_1 + x_2}{2} = x_{3}\] , where x3 is the x-component of the midpoint

OpenStudy (ipwnbunnies):

\[\frac{3 + x_{2}}{2} = 5\] Solve for x2, that'll be the x-value of the remaining end point. Do the same for the y-value.

OpenStudy (anonymous):

Okay, thank you @iPwnBunnies , but what y value?

OpenStudy (anonymous):

I know of the endpoint, but where would I plug it in?

OpenStudy (ipwnbunnies):

It's just another point you're trying to find. So, like when you solve for x2, you get 7. (7, Y) will be the endpoint you're trying to find. To find the y-value, use the same formula, and the use y-values from the given endpoint and midpoint. \[\frac{5 + y_{2}}{2} = -11\]

OpenStudy (anonymous):

Ohh, thank you :) @iPwnBunnies

OpenStudy (anonymous):

@iPwnBunnies I got 3i , is this right?